# Spin transport and lack of quantisation in the $A\mathrm{II}$ class on the honeycomb structure

### v1

@@ 3.4. Spin-non-conserving contributions
Let us first recall what happens when the spin $S_{z}$ is conserved, that is, when $[H,S_{z}]=0$. For any orthogonal projection $P\in\mathcal{P}(\mathcal{H})$, one defines its (first) Chern number by
@@
[…]
@@
Let $(H,\mu)$ be an insulator with $H\in\mathcal{P}_{0}(\mathcal{H})\cap\mathrm{AII}^{\hexagon}$ such that $[H,S_{z}]=0$. Then,
@@
In other words, whenever the spin $S_{z}$ is conserved, the spin conductivity reduces to two copies of the charge conductivity, associated with the different spin sectors, and so its quantisation in fundamental units $\frac{e}{h}=\frac{1}{2\pi}$ is obviously established.
@@
For brevity, we write $P^{\upuparrows}:=p^{\uparrow}Pp^{\uparrow}$ and $P^{\downdownarrows}:=p^{\downarrow}Pp^{\downarrow}$, so that $[H,S_{z}]=0$ implies $P=P^{\upuparrows}+P^{\downdownarrows}$. Thus, by using Proposition 3.13(iii), Lemma 3.2(i) and (iii) one obtains that
$\begin{split}\sigma_{ij}^{\mathrm{s}}&=\operatorname{Re}\tau\big{(}\mathrm{i}S_{z}P[[X_{i},P],[X_{j},P]]\big{)}\\ &=\frac{1}{2}\tau\big{(}\mathrm{i}P^{\upuparrows}[[X_{i},P^{\upuparrows}],[X_{j},P^{\upuparrows}]]\big{)}-\frac{1}{2}\tau\big{(}\mathrm{i}P^{\downdownarrows}[[X_{i},P^{\downdownarrows}],[X_{j},P^{\downdownarrows}]]\big{)}\\ &=\frac{1}{2\pi}\left(\frac{1}{2}\mbox{Chern}(P^{\upuparrows})_{ij}-\frac{1}{2}\mbox{Chern}(P^{\downdownarrows})_{ij}\right)=\frac{1}{2\pi}\mbox{ $S$-Chern$(P)_{ij}$}\;.\end{split}$
where we used that the fibre of $[X_{i},P]$ is $\mathrm{i}\partial_{k_{i}}P(k)$. To prove that this is valued in $\mathbb{Z}/2\pi$, we note the identity $\mbox{Chern}(P^{\downdownarrows})=-\mbox{Chern}(P^{\upuparrows})$ due to the time-reversal symmetry, and conclude that $\sigma_{ij}^{\mathrm{s}}=\frac{1}{2\pi}\mbox{Chern}(P^{\upuparrows})_{ij}\in\frac{1}{2\pi}\mathbb{Z}$. ∎
In the general case in which the spin $S_{z}$ is not conserved, that is, when $[H,S_{z}]\neq 0$, as anticipated in the Introduction, we decompose the Hamiltonian into its spin-commuting term $H^{\mathrm{sc}}$ and spin-non-commuting term $H^{\mathrm{snc}}$, which we here recall:
@@
${\left|\kern-1.07639pt\left|\kern-1.07639pt\left|A\right|\kern-1.07639pt\right|\kern-1.07639pt\right|}:=\left\|A\right\|+\sum_{j=1}^{2}\left\|[A,X_{j}]\right\|,$ (3.30)
whenever the commutator $[A,X_{j}]$ (is densely defined and) extends to a bounded operator.
@@
We say that an insulator $(H,\mu)$ almost conserves the spin (in the $z$-direction) if and only if there exists a sufficiently small constant $C_{s}$ such that ${\left|\kern-1.07639pt\left|\kern-1.07639pt\left|[H,S_{z}]\right|\kern-1.07639pt\right|\kern-1.07639pt\right|}\leq C_{s}$, or equivalently ${\left|\kern-1.07639pt\left|\kern-1.07639pt\left|H^{\mathrm{snc}}\right|\kern-1.07639pt\right|\kern-1.07639pt\right|}\leq C_{s}$.
@@
    Notice that we require that $C_{s}$ is smaller than the size of the gap associated with the insulator $(H,\mu)$ in order to have that $(H^{\mathrm{sc}},\mu)$ is an insulator as well (thus the scalar $C_{s}$ depends on $H$ and $\mu$). Indeed, let $z$ be in the spectral gap of $H$ then
    $H^{\mathrm{sc}}-z$ $=((H^{\mathrm{sc}}-H){(H-z)}^{-1}+\mathbb{1})(H-z)=(\mathbb{1}-H^{\mathrm{snc}}{(H-z)}^{-1})(H-z),$
    the right-hand side operator is invertible if $\left\|H^{\mathrm{snc}}\right\|\left\|{(H-z)}^{-1}\right\|\leq C_{s}\left\|{(H-z)}^{-1}\right\|<1$. In this case, we can define the Fermi projector $P^{\mathrm{sc}}:=P(H^{\mathrm{sc}},\mu)$ associated with $(H^{\mathrm{sc}},\mu)$ by the Riesz formula:
    $P^{\mathrm{sc}}=\frac{\mathrm{i}}{2\pi}\oint_{\mathcal{C}}\mathrm{d}z\,(H^{\mathrm{sc}}-z)^{-1},$ (3.31)
@@
For insulators in the $\mathrm{AII}^{\hexagon}$ class that almost conserve the spin, we derive approximation results for the spin conductivity $\sigma^{\mathrm{s}}_{ij}$ using an approximation procedure [Pr, Sch, MaPaTe]. In the following theorem, which is the main result of this section, we will show that $\sigma^{\mathrm{s}}_{ij}$ equals, up to a remainder of order ${\left|\kern-1.07639pt\left|\kern-1.07639pt\left|H^{\mathrm{snc}}\right|\kern-1.07639pt\right|\kern-1.07639pt\right|}^{2}$, both the spin-commuting term $\Sigma^{\mathrm{sc}}_{ij}$ in (3.19) and the $\frac{1}{2\pi}$-spin Chern number $\frac{1}{2\pi}\mbox{$S$-Chern}(P^{\mathrm{sc}})_{ij}$ associated with the spin-commuting Fermi projection, see (3.26) and Remark 3.21. These results refine those obtained in [Sch, Eq. (8)] and [MaPaTe, Proposition 5.13], improving them by one order in ${\left|\kern-1.07639pt\left|\kern-1.07639pt\left|[H,S_{z}]\right|\kern-1.07639pt\right|\kern-1.07639pt\right|}$.
@@
Let $(H,\mu)$ be an insulator with $H\in\mathcal{P}_{0}(\mathcal{H})\cap\mathrm{AII}^{\hexagon}$ such that it almost conserves the spin in the sense of Definition 3.20. Let $P=P(H,\mu)$ be its Fermi projection. Then there exists a constant $C$ independent of ${\left|\kern-1.07639pt\left|\kern-1.07639pt\left|[H,S_{z}]\right|\kern-1.07639pt\right|\kern-1.07639pt\right|}$ such that
$\big{|}\sigma^{\mathrm{s}}_{ij}-\Sigma^{\mathrm{sc}}_{ij}\big{|}\leq C{\left|\kern-1.07639pt\left|\kern-1.07639pt\left|[H,S_{z}]\right|\kern-1.07639pt\right|\kern-1.07639pt\right|}^{2}$ (3.32)
@@
$\big{|}\sigma^{\mathrm{s}}_{ij}-\frac{1}{2\pi}\mbox{$S$-Chern}(P^{\mathrm{sc}})_{ij}\big{|}\leq C{\left|\kern-1.07639pt\left|\kern-1.07639pt\left|[H,S_{z}]\right|\kern-1.07639pt\right|\kern-1.07639pt\right|}^{2}\;,$ (3.33)
@@
Note that the deviation (3.32) was numerically studied in [MoUl] for the Kane–Mele and for the Bernevig–Hughes–Zhang [BHZ] models, shown to be quadratic in ${\left|\kern-1.07639pt\left|\kern-1.07639pt\left|[H,S_{z}]\right|\kern-1.07639pt\right|\kern-1.07639pt\right|}$, in agreement with our analytic result.
@@
$\mathcal{Z}(S_{z}):=\left\{A\in\mathcal{P}(\mathcal{H}):\;\exists\;C>0\text{ such that }\left\|[A,S_{z}]\right\|\leq C{\left|\kern-1.07639pt\left|\kern-1.07639pt\left|[H,S_{z}]\right|\kern-1.07639pt\right|\kern-1.07639pt\right|}\right\},$ (3.34)
bearing in mind that we are interested in models, like the extended Kane–Mele model in Subsection 4.1, where ${\left|\kern-1.07639pt\left|\kern-1.07639pt\left|[H,S_{z}]\right|\kern-1.07639pt\right|\kern-1.07639pt\right|}$ is a small quantity, equivalently $H^{\mathrm{snc}}$ is a small perturbation of $H^{\mathrm{sc}}$, see (3.28). In other words, $\mathcal{Z}(S_{z})$ consists of all periodic bounded operators which “nearly commute” with the spin operator.
@@
Let $(H,\mu)$ be an insulator that almost conserves the spin. If $A,B,F\in\mathcal{Z}(S_{z})$, then there exists a constant $C$ independent of ${\left|\kern-1.07639pt\left|\kern-1.07639pt\left|[H,S_{z}]\right|\kern-1.07639pt\right|\kern-1.07639pt\right|}$ such that
$\left\lvert\tau\left(A[S_{z},F]B\right)\right\rvert\leq C{\left|\kern-1.07639pt\left|\kern-1.07639pt\left|[H,S_{z}]\right|\kern-1.07639pt\right|\kern-1.07639pt\right|}^{2}\;.$ (3.35)
@@
$[P,S_{z}]=\frac{\mathrm{i}}{2\pi}\oint_{\mathcal{C}}\mathrm{d}z\,[(H-z)^{-1},S_{z}]\in\mathcal{Z}(S_{z}).$ (3.39)
@@
$[[H,X_{j}],S_{z}]=[[H,S_{z}],X_{j}],$
so we obtain that $[H,X_{j}]\in\mathcal{Z}(S_{z})$ as well. By Jacobi’s identity, (3.38) and (3.37), we have that
@@
thus one estimates the right-hand side terms by ${\left|\kern-1.07639pt\left|\kern-1.07639pt\left|[H,S_{z}]\right|\kern-1.07639pt\right|\kern-1.07639pt\right|}$ up to a constant, being uniform in $z\in\mathcal{C}$. Thus, again in view of (3.6), we get that
$[[P,X_{j}],S_{z}]=\frac{\mathrm{i}}{2\pi}\oint_{\mathcal{C}}\mathrm{d}z\,[[(H-z)^{-1},X_{j}],S_{z}]\in\mathcal{Z}(S_{z}).$
@@
$[P_{1},S_{z}]=\frac{\mathrm{i}}{2\pi}\oint_{\mathcal{C}}\mathrm{d}z\,[(H-z)^{-1}[[X_{j},P],P](H-z)^{-1},S_{z}],$
@@
By Lemma 3.25 and Remark 3.23 $[H,X_{i}],P,PP_{1},[H,X_{i}]P,P_{1}\in\mathcal{Z}(S_{z})$, so Lemma 3.24 implies that
$\big{|}\tau\left(\mathrm{i}[H,X_{i}][[S_{z},P],P]P_{1}\right)\big{|}\leq C{\left|\kern-1.07639pt\left|\kern-1.07639pt\left|[H,S_{z}]\right|\kern-1.07639pt\right|\kern-1.07639pt\right|}^{2}\;.$
@@
Observe that the operator $S_{z}(H^{\mathrm{sc}}-z)^{-1}[[X_{i},P],[X_{j},P]]$ is in $\mathcal{Z}(S_{z})$ in view of Lemma 3.25 and Remark 3.23. Also the operators $H,S_{z}(H-z)^{-1}\in\mathcal{Z}(S_{z})$ again by Lemma 3.25, thus Lemma 3.24 implies that the first summand on the r.h.s. of (3.41) is bounded, up to a constant, by ${{\left|\kern-1.07639pt\left|\kern-1.07639pt\left|[H,S_{z}]\right|\kern-1.07639pt\right|\kern-1.07639pt\right|}}^{2}$. For the analysis of the second term on the r.h.s. of (3.41), we note that by Leibniz’s rule and (3.42), we get that
@@
.$\phantom{=}\Bigl{.}+(H-z)^{-1}][[H,X_{i}],S_{z}]S_{z}(H^{\mathrm{sc}}-z)^{-1}+(H-z)^{-1}][S_{z},H]S_{z}[X_{i},(H^{\mathrm{sc}}-z)^{-1}]\Bigr{)}$
@@
.$\phantom{=}\Bigl{.}-\tau\left(\mathrm{i}S_{z}P^{\mathrm{sc}}[X_{j},P][X_{i},(H-z)^{-1}][S_{z},H]S_{z}(H^{\mathrm{sc}}-z)^{-1}\right)\Bigr{)}.$
Since on the r.h.s of the last equality the operators $S_{z}P^{\mathrm{sc}}[X_{i},(H-z)^{-1}],\,H,\,S_{z}(H^{\mathrm{sc}}-z)^{-1}[X_{j},P]\in\mathcal{Z}(S_{z})$ and also the operators $S_{z}P^{\mathrm{sc}}[X_{j},P][X_{i},(H-z)^{-1}],\,S_{z}(H^{\mathrm{sc}}-z)^{-1}\in\mathcal{Z}(S_{z})$ by using again Lemma 3.25 and Remark 3.23, then Lemma 3.24 implies that the l.h.s. of (3.43) is bounded up to a constant by ${\left|\kern-1.07639pt\left|\kern-1.07639pt\left|[H,S_{z}]\right|\kern-1.07639pt\right|\kern-1.07639pt\right|}^{2}$. The analysis of the contributions for the second term on the r.h.s. of (3.41) coming from $T_{2}$ and $T_{3}$ is analogous. The last summand on the r.h.s. of (3.41) can be estimated similarly. ∎
@@ 4.1. The extended Kane–Mele model
We here consider a simple generalisation of the Kane–Mele model in which a next-to-nearest-neighbour Rashba interaction, being compatible with the $\mathrm{AII}^{\hexagon}$ class symmetries, is added. This additional term is denoted below by $H_{\mathrm{R2}}$. This model, which we refer to as extended Kane–Mele model, is characterised by the following Hamilton operator $H_{\mathrm{KM}}=H_{\mathrm{KM}}(t,\lambda_{\mathrm{SO}},w,\lambda_{\mathrm{R}},r)$
@@
[…]
@@
[…]
@@
Note that $H_{\mathrm{KM}}$ is not spin-conserving, that is $[H_{\mathrm{KM}},S_{z}]\neq 0$ because of the Rashba Hamiltonian, indeed:
$[H-H_{\mathrm{R}},S_{z}]=0\quad\text{and}\quad[H_{\mathrm{R}},S_{z}]\neq 0\;.$
@@
For any $t,\lambda_{\mathrm{SO}},w,\lambda_{\mathrm{R}},\lambda_{\mathrm{R2}}\in\mathbb{R}$, we have that $H_{\mathrm{KM}}\in\mathrm{AII}^{\hexagon},$ where the $\mathrm{AII}^{\hexagon}$ class is specified as in Definition 2.4.
@@
Note that the operators $\chi_{A}$, $\chi_{B}$, and $T_{v}$ for any fixed $v\in\mathbb{R}^{2}$, are short-range, so is $H_{\mathrm{KM}}$. Furthermore, $H_{\mathrm{KM}}$ is periodic since $[T_{v},T_{u}]=0$ for any vectors $v,u\in\mathbb{R}^{2}$ and $[\chi_{\sharp},T_{\gamma}]=0$ for every $\sharp\in\{A,B\}$.
To verify that $H_{\mathrm{KM}}$ is $32\pi/$ 3-rotationally symmetric, one notes that
@@
Analogously, one checks that $R_{2\pi/3}H_{\mathrm{R2}}=H_{\mathrm{R2}}R_{2\pi/3}$. Using again (4.3), we obtain $[R_{2\pi/3},H_{\mathrm{NN}}]=[R_{2\pi/3},H_{\mathrm{SO}}]=0$, and obviously $[R_{2\pi/3},H_{v}]=0$ as well. Hence, $H_{\mathrm{KM}}$ is $32\pi/$ 3-rotationally symmetric.
Let us proceed to check the invariance of $H_{\mathrm{KM}}$ under $\mathrm{v\otimes h}$-reflection. By (4.2), we have
@@
Furthermore, we observe that the $\mathrm{v\otimes h}$-reflection and translation operators are intertwined by the following identity:
$R_{\mathrm{v\otimes h}}\,T_{v}=T_{\mathrm{R}_{\mathrm{v}}v}\,R_{\mathrm{v\otimes h}}.$ (4.5)
By identity (4.5), one easily verifies that $[H_{\mathrm{NN}},R_{\mathrm{v\otimes h}}]=[H_{v},R_{\mathrm{v\otimes h}}]=0$. By using the third equality in (4.4) and (4.5), one has that $[H_{\mathrm{SO}},R_{\mathrm{v\otimes h}}]=0$. By employing the first two equalities in (4.4) and again (4.5), one has also that
$3R_{\mathrm{v\otimes h}}H_{\mathrm{R1}}$ $=\mathrm{i}R_{\mathrm{v\otimes h}}\Big{[}-\big{(}T_{d_{1}}-T_{-d_{1}}\big{)}\otimes\sigma_{+}+\big{(}T_{d_{2}}-T_{-d_{2}}\big{)}\otimes\sigma_{-}+\big{(}T_{d_{3}}-T_{-d_{3}}\big{)}\otimes\sigma_{y}\Big{]}$
$=\mathrm{i}\Big{[}\big{(}T_{d_{2}}-T_{-d_{2}}\big{)}\otimes\sigma_{-}+\big{(}T_{d_{1}}-T_{-d_{1}}\big{)}\otimes\left(-\sigma_{+}\right)+\big{(}T_{d_{3}}-T_{-d_{3}}\big{)}\otimes\sigma_{y}\Big{]}R_{\mathrm{v\otimes h}}$
$=3H_{\mathrm{R1}}R_{\mathrm{v\otimes h}}.$
Similarly, one checks that $[H_{\mathrm{R2}},R_{\mathrm{v\otimes h}}]=0$. By similar computations, one also verifies that $[H_{\mathrm{KM}},\Theta]=0$ and that $\{H_{\mathrm{R}},\Xi\}\neq 0$. ∎

### v2

@@ 3.4. Spin-non-conserving contributions
Let us first recall what happens when the spin $S_{z}$ is conserved, that is, when $H,S_{z}]=0$. For any orthogonal projection $P\in\mathcal{P}(\mathcal{H})$, one defines its (first) Chern number by
@@
 $\mbox{ $S$-Chern$(P)_{ij}$}:=\frac{1}{2}\mbox{Chern}(p^{\uparrow}Pp^{\uparrow})_{ij}-\frac{1}{2}\mbox{Chern}(p^{\downarrow}Pp^{\downarrow})_{ij}\in\frac{1}{2}\mathbb{Z}\;.$ (3.26)
@@
Let $(H,\mu)$ be an insulator with $H\in\mathcal{P}_{0}^{\hexagon}(\mathcal{H})$ such that $H,S_{z}]=0$. Then,
$\sigma_{ij}^{\mathrm{s}}=\frac{1}{2\pi}\mbox{ $S$-Chern$(P)_{ij}$}\;.$
Furthermore, if $H$ is time-reversal symmetric, see (2.15) and below, then
@@
In other words, whenever the spin $S_{z}$ is conserved and $H$ is time-reversal symmetric, the spin conductivity reduces to two copies of the charge conductivity, associated with the different spin sectors, and so its quantisation in fundamental units $\frac{e}{h}=\frac{1}{2\pi}$ is obviously established.
@@
For brevity, we write $P^{\upuparrows}:=p^{\uparrow}Pp^{\uparrow}$ and $P^{\downdownarrows}:=p^{\downarrow}Pp^{\downarrow}$, so that $H,S_{z}]=0$ implies $P=P^{\upuparrows}+P^{\downdownarrows}$. Thus, by using Proposition 3.13(iii), Lemma 3.2(i) and (iii) one obtains that
$\begin{split}\sigma_{ij}^{\mathrm{s}}&=\operatorname{Re}\tau\big(\mathrm{i}S_{z}P[[X_{i},P],[X_{j},P]]\big)\\ &=\frac{1}{2}\tau\big(\mathrm{i}P^{\upuparrows}[[X_{i},P^{\upuparrows}],[X_{j},P^{\upuparrows}]]\big)-\frac{1}{2}\tau\big(\mathrm{i}P^{\downdownarrows}[[X_{i},P^{\downdownarrows}],[X_{j},P^{\downdownarrows}]]\big)\\ &=\frac{1}{2\pi}\left(\frac{1}{2}\mbox{Chern}(P^{\upuparrows})_{ij}-\frac{1}{2}\mbox{Chern}(P^{\downdownarrows})_{ij}\right)=\frac{1}{2\pi}\mbox{ $S$-Chern$(P)_{ij}$}\in\frac{1}{4\pi}\mathbb{Z}\;,\end{split}$
where we used that the fibre of $X_{i},P]$ is $\mathrm{i}\partial_{k_{i}}P(k)$. Moreover, if $H$ is time-reversal symmetric then we note the identity $\mbox{Chern}(P^{\downdownarrows})=-\mbox{Chern}(P^{\upuparrows})$ and conclude that $\sigma_{ij}^{\mathrm{s}}=\frac{1}{2\pi}\mbox{Chern}(P^{\upuparrows})_{ij}\in\frac{1}{2\pi}\mathbb{Z}$. ∎
In the general case in which the spin $S_{z}$ is not conserved, that is, when $H,S_{z}]\neq 0$, as anticipated in the Introduction, we decompose the Hamiltonian into its spin-commuting term $H^{\mathrm{sc}}$ and spin-non-commuting term $H^{\mathrm{snc}}$, which we here recall:
@@
[…]
whenever the commutator $A,X_{j}]$ (is densely defined and) extends to a bounded operator.
@@
[…]
@@
    Note that requiring that $C_{s}$ is smaller than the size of the gap associated with the insulator $(H,\mu)$ implies that $(H^{\mathrm{sc}},\mu)$ is an insulator as well. Indeed, let $\mathcal{C}$ be the complex contour in (3.6), fix $z\in\mathcal{C}$, consider
    $H^{\mathrm{sc}}-z$ $=((H^{\mathrm{sc}}-H){(H-z)}^{-1}+\mathbb{1})(H-z)=(\mathbb{1}-H^{\mathrm{snc}}{(H-z)}^{-1})(H-z).$
    Since $\left\|H^{\mathrm{snc}}\right\|\left\|{(H-z)}^{-1}\right\|\leq\frac{C_{s}}{\operatorname{dist}(z,\operatorname{Spectrum}(H))}\leq\frac{C_{s}}{\operatorname{dist}(\mu,\operatorname{Spectrum}(H))}<1$, the operator on the right-hand side is invertible. Thus, we can define the Fermi projector $P^{\mathrm{sc}}:=P(H^{\mathrm{sc}},\mu)$ associated with $(H^{\mathrm{sc}},\mu)$ by the Riesz formula:
    $P^{\mathrm{sc}}=\frac{\mathrm{i}}{2\pi}\oint_{\mathcal{C}}\mathrm{d}z\,(H^{\mathrm{sc}}-z)^{-1},$
@@
[…]
@@
[…]
@@
[…]
@@
Under the hypotheses of Theorem 3.22 and assuming that $H$ is time-reversal symmetric, then Theorem 1.1(ii) follows.
@@
[…]
@@
[…]
@@
$P,S_{z}]=\frac{\mathrm{i}}{2\pi}\oint_{\mathcal{C}}\mathrm{d}z\,[(H-z)^{-1},S_{z}]\in\mathcal{Z}(S_{z}).$
@@
$H,X_{j}],S_{z}]=[[H,S_{z}],X_{j}],$
so we obtain that $H,X_{j}]\in\mathcal{Z}(S_{z})$ as well. By Jacobi’s identity, (3.37) and (3.36), we have that
@@
[…]
$P,X_{j}],S_{z}]=\frac{\mathrm{i}}{2\pi}\oint_{\mathcal{C}}\mathrm{d}z\,[[(H-z)^{-1},X_{j}],S_{z}]\in\mathcal{Z}(S_{z}).$
@@
$P_{1},S_{z}]=\frac{\mathrm{i}}{2\pi}\oint_{\mathcal{C}}\mathrm{d}z\,[(H-z)^{-1}[[X_{j},P],P](H-z)^{-1},S_{z}],$
@@
By Lemma 3.25 and Remark 3.23 $H,X_{i}],P,PP_{1},[H,X_{i}]P,P_{1}\in\mathcal{Z}(S_{z})$, so Lemma 3.24 implies that
[…]
@@
[…]
@@
$\phantom{=}\Bigl.+(H-z)^{-1}][[H,X_{i}],S_{z}]S_{z}(H^{\mathrm{sc}}-z)^{-1}+(H-z)^{-1}][S_{z},H]S_{z}[X_{i},(H^{\mathrm{sc}}-z)^{-1}]\Bigr)$
@@
$\phantom{=}\Bigl.-\tau\left(\mathrm{i}S_{z}P^{\mathrm{sc}}[X_{j},P][X_{i},(H-z)^{-1}][S_{z},H]S_{z}(H^{\mathrm{sc}}-z)^{-1}\right)\Bigr).$
[…]
@@ 4.1. The extended Kane–Mele model
We consider a simple generalisation of the Kane–Mele model obtained by adding a next-to-nearest-neighbour Rashba interaction compatible with its symmetries, denoted by $H_{\mathrm{R2}}$. The resulting extended Kane–Mele model is characterised by the Hamiltonian $H_{\mathrm{KM}}=H_{\mathrm{KM}}(t,\lambda_{\mathrm{SO}},w,\lambda_{\mathrm{R}},r)$
@@
where $t>0,\lambda_{\mathrm{SO}}\in\mathbb{R}\setminus\{0\}$, $w,\lambda_{\mathrm{R}},r\in\mathbb{R}$ are strength parameters (6) (6) (6)Note that both $t$ and $\lambda_{\mathrm{SO}}$ are assumed to be non-zero otherwise the physics of the model is completely different. Moreover, observe that our choice of parameters reflects the fact that $H_{\mathrm{R2}}$ is a second-order effect with respect to $H_{\mathrm{R1}}$, so that if $H_{\mathrm{R1}}$ is absent, so is $H_{\mathrm{R2}}$., and where $H_{\mathrm{\sharp}}$ for $\sharp\in\{\mathrm{NN},\mathrm{SO},\mathrm{W},\mathrm{R1},\mathrm{R2}\}$ are specified below, as operators acting on $\ell^{2}(\mathcal{C})\otimes\mathbb{C}^{2}$:
@@
The numerical prefactors $\frac{1}{3\sqrt{3}}$ and $\frac{1}{3}$ that appear in the definition of $H_{\mathrm{\sharp}}$, for $\sharp\in\{\mathrm{SO},\mathrm{R1},\mathrm{R2}\}$ are there only for convenience of the subsequent analysis.
@@
Note that $H_{\mathrm{KM}}$ is not spin-conserving, that is $H_{\mathrm{KM}},S_{z}]\neq 0$ because of the Rashba Hamiltonian, indeed:
$H-H_{\mathrm{R}},S_{z}]=0\quad\text{and}\quad[H_{\mathrm{R}},S_{z}]\neq 0\;.$
@@
For any $t,\lambda_{\mathrm{SO}},w,\lambda_{\mathrm{R}},r\in\mathbb{R}$, the Hamiltonian $H_{\mathrm{KM}}$ belongs to $\mathcal{P}_{0}^{\hexagon}(\mathcal{H})$ and is inversion and time-reversal symmetric, compare with (2.14) and (2.15).
@@
Note that the operators $\chi_{A}$, $\chi_{B}$, and $T_{v}$ for any fixed $v\in\mathbb{R}^{2}$, are short-range, so is $H_{\mathrm{KM}}$. Furthermore, $H_{\mathrm{KM}}$ is periodic since $T_{v},T_{u}]=0$ for any vectors $v,u\in\mathbb{R}^{2}$ and $\chi_{\sharp},T_{\gamma}]=0$ for every $\sharp\in\{A,B\}$.
To verify that $H_{\mathrm{KM}}$ is $2\pi/3$-rotationally symmetric, one notes that
@@
Analogously, one checks that $R_{2\pi/3}H_{\mathrm{R2}}=H_{\mathrm{R2}}R_{2\pi/3}$. Using again (4.3), we obtain $R_{2\pi/3},H_{\mathrm{NN}}]=[R_{2\pi/3},H_{\mathrm{SO}}]=0$, and obviously $R_{2\pi/3},H_{v}]=0$ as well. Hence, $H_{\mathrm{KM}}$ is $2\pi/3$-rotation symmetric.
Let us proceed to check that $H_{\mathrm{KM}}$ is inversion symmetric. The relevant operator in this case is
$(\Pi_{1}\psi)(x):=\mathrm{e}^{-\mathrm{i}\pi S_{y}}\psi(\mathrm{R}_{v}^{-1}x)\;,\qquad\forall\psi\in\mathcal{H}\;,$
where $\mathrm{R}_{v}$ is the restriction to $\mathcal{C}$ of the vertical reflection on $\mathbb{R}^{2}$, namely $\mathrm{R}_{v}(x_{1},x_{x})=(x_{1},-x_{2})$. It is straightforward to check that $\Pi_{1}$ satisfies (2.14) with $a=1$. By (4.2), we have
@@
Furthermore, we observe that the $\Pi_{1}$ and translation operators are intertwined by the following identity:
$\Pi_{1}\,T_{v}=T_{\mathrm{R}_{\mathrm{v}}v}\,\Pi_{1}.$ (4.5)
By identity (4.5), one easily verifies that $H_{\mathrm{NN}},\Pi_{1}]=[H_{v},\Pi_{1}]=0$. By using the third equality in (4.4) and (4.5), one has that $H_{\mathrm{SO}},\Pi_{1}]=0$. By employing the first two equalities in (4.4) and again (4.5), one also has that
$3\Pi_{1}H_{\mathrm{R1}}$ $=\mathrm{i}\Pi_{1}\Big[-\big(T_{d_{1}}-T_{-d_{1}}\big)\otimes\sigma_{+}+\big(T_{d_{2}}-T_{-d_{2}}\big)\otimes\sigma_{-}+\big(T_{d_{3}}-T_{-d_{3}}\big)\otimes\sigma_{y}\Big]$
$=\mathrm{i}\Big[\big(T_{d_{2}}-T_{-d_{2}}\big)\otimes\sigma_{-}+\big(T_{d_{1}}-T_{-d_{1}}\big)\otimes\left(-\sigma_{+}\right)+\big(T_{d_{3}}-T_{-d_{3}}\big)\otimes\sigma_{y}\Big]\Pi_{1}$
$=3H_{\mathrm{R1}}\Pi_{1}.$
Similarly, one checks that $H_{\mathrm{R2}},\Pi_{1}]=0$. Finally, analogous computations show that $H_{\mathrm{KM}},\Theta]=0$. ∎